The weight of a piece of metal is 225 lbf at a location where gravitational acceleration is 31 ft/s2 . Find the weight of the metal in lbf and the mass of the metal in lbm if it will be placed on the surface of the moon where g is 5.5 ft/s2 .


1
Expert's answer
2021-09-09T00:33:38-0400

"F = m\times a"

"m =\frac{F}{a} =\frac{225}{31} = 7.25 lbm"

mass of the metal = 7.25 lbm

Weight of metal

"W =m \times g = 7.25 \times 5.5 = 39.875" lbf